# A Unified Voxel Diffusion Module for Point Cloud 3D Object Detection

### v1

@@ Abstract
Recent advances in point cloud object detection have increasingly adopted Transformer-based and State Space Models (SSMs), demonstrating strong performance. However, voxel-based representations in these models require strict consistency in input and output dimensions due to their serialized processing, which limits the spatial diffusion capability typically offered by convolutional operations. This limitation significantly affects detection accuracy. Inspired by CNN-based object detection architectures, we propose a novel Voxel Diffusion Module (VDM) to enhance voxel-level representation and diffusion in point cloud data. VDM is composed of sparse 3D convolutions, submanifold sparse convolutions, and residual connections. To ensure computational efficiency, the output feature maps are downsampled to one-fourth of the original input resolution. VDM serves two primary functions: (1) diffusing foreground voxel features through sparse 3D convolutions to enrich spatial context, and (2) aggregating fine-grained spatial information to strengthen voxel-wise feature representation. The enhanced voxel features produced by VDM can be seamlessly integrated into mainstream Transformer- or SSM-based detection models for accurate object classification and localization, highlighting the generalizability of our method. We evaluate VDM on several benchmark datasets by embedding it into both Transformer-based and SSM-based models. Experimental results show that our approach consistently improves detection accuracy over baseline models. Specifically, VDM-SSMs achieve 74.7 mAPH (L2) on Waymo, 72.9 NDS on nuScenes, 42.3 mAP on Argoverse 2, and 67.6 mAP on ONCE, setting new state-of-the-art performance across all datasets. Our code will be made publicly available.
Introduction
@@
3D object detection based on point clouds plays a crucial role in applications such as autonomous driving (Bansal, Krizhevsky, and Ogale 2018) and robotic navigation (Zhu et al. 2017). Inspired by the remarkable success of Transformers (Vaswani et al. 2017) and State Space Models (SSMs) (Gu and Dao 2023) in natural language processing, researchers have recently explored adapting these architectures to computer vision tasks, achieving strong performance. For example, in the domain of point cloud 3D object detection, DVST (Wang et al. 2023) introduces a dynamic sparse window attention mechanism that adaptively partitions each window into local sub-regions based on voxel sparsity, enabling fully parallel attention computations. LION (Liu et al. 2024) further extends this direction by employing large grouping windows and leveraging SSMs’ ability to capture long-range voxel dependencies, setting new state-of-the-art (SOTA) performance on several public datasets.
Figure 1: Performance comparison of different models on various datasets. Numbers highlighted in red indicate that our model achieves state-of-the-art results.
Despite these advancements, both Transformer-based and SSM-based models for point cloud voxel sequence processing require consistency in input and output sequence dimensions, which fundamentally limits their ability to perform convolution-like spatial diffusion. However, prior work (Zhang et al. 2024a; Liu et al. 2024) has demonstrated that voxel diffusion, particularly for foreground voxels, is critical for enhancing detection accuracy. For instance, LION selects top-$k$ voxels via feature summation and applies directional diffusion during intermediate processing. In contrast, our method proposes to conduct voxel diffusion before serialization, using convolutional operations to enrich the original voxel space with contextually relevant features.
Motivated by CNN-based object detection architectures, we propose a novel Voxel Diffusion Module (VDM) tailored for both Mamba-based and Transformer-based 3D detection models. Positioned before voxel sequence serialization, VDM leverages 3D convolutional operations to diffuse foreground voxel features and increase their spatial density. Simultaneously, sparse convolutions are employed to aggregate fine-grained spatial features within each voxel’s local 3D neighborhood, supplying the downstream sequence model with richer spatial context. Structurally, VDM consists of stacked sparse 3D convolutions and residual blocks, effectively supporting both diffusion and aggregation. The output feature map is downsampled to one-fourth of the input resolution (in height and width), helping reduce the computational load for subsequent models.
@@
VDM is designed to be modular and general, allowing seamless integration with both Transformer-based and SSM-based detectors. To evaluate its effectiveness, we perform extensive experiments on four large-scale datasets. As shown in Figure 1, our VDM-enhanced SSM-based model (VDM-Mamba) achieves 74.7 mAPH (L2) on Waymo, 68.1 mAP on nuScenes, 42.3 mAP on Argoverse 2, and 67.6 mAP on ONCE, setting a new state-of-the-art (SOTA) performance. Additionally, the Transformer-based variant (VDM-DSVT) improves performance by +2.2 mAPH (L2) over the DSVT baseline on the Waymo dataset, confirming the general effectiveness of our approach.
@@
  • We propose VDM, a voxel diffusion module that enhances the representation of raw voxel inputs through 3D convolution, significantly enriching foreground voxel features and improving detection accuracy.
  • VDM utilizes sparse 3D convolutions and residual networks to aggregate spatial features from the voxel neighborhood, providing richer positional cues for downstream sequence modeling.
  • VDM is a general-purpose module that integrates well with both Transformer-based and SSM-based 3D detectors, consistently boosting model performance.
  • We validate the effectiveness of VDM through experiments on multiple public datasets, where VDM-SSMs achieve SOTA results, and VDM also improves Transformer-based models.
Related Work
3D object detection based on point clouds has seen significant progress in recent years. According to model architecture, existing approaches can be broadly categorized into three types: CNN-based, Transformer-based, and SSM-based methods. CNN-based models emerged earlier and have been thoroughly studied, achieving solid performance. In contrast, Transformer and Mamba-based models are relatively recent, and while less explored for 3D object detection, they have already shown promising results—often surpassing CNN-based methods in detection accuracy.
CNN-Based 3D Object Detection. CNN-based 3D detectors are typically categorized into voxel-based and point-based approaches. VoxelNet (Zhou and Tuzel 2018) introduces a Voxel Feature Encoding (VFE) layer to convert point clusters within voxels into unified feature representations. HEDNet uses an encoder-decoder design to capture long-range spatial dependencies, particularly useful for detecting large or distant objects. SAFDNet proposes an adaptive feature diffusion strategy to address missing center features in sparse detectors. PointPillars (Lang et al. 2019) leverages PointNet to learn features from vertical columns of point clouds, offering notable inference speed advantages. CenterPoint (Yin, Zhou, and Krahenbuhl 2021) detects object centers using a keypoint detector and regresses attributes such as 3D size, orientation, and velocity. 3DSSD (Yang et al. 2020) introduces a fused sampling strategy in the downsampling process to retain informative representative points. IASSD (Zhang et al. 2022) proposes two task-aware, instance-sensitive sampling strategies for hierarchically selecting foreground points of interest. Among these approaches, voxel-based detectors generally achieve higher accuracy and have become the dominant choice for outdoor 3D detection tasks.
Transformer-Based 3D Object Detection. Transformers have achieved state-of-the-art performance in both NLP and computer vision tasks. PVT-SSD (Yang et al. 2023) introduces a Point-Voxel Transformer that captures long-range dependencies via voxel tokens while preserving fine spatial detail through point-level features. FlatFormer (Liu et al. 2023) flattens point clouds using window-based sorting and partitions them into equal-sized groups, improving detection efficiency. DSVT (Wang et al. 2023) proposes a Dynamic Sparse Window Attention mechanism that adaptively splits each window based on voxel sparsity, enabling fully parallel feature computation. However, the serialization of voxel features in Transformer-based methods can lead to the loss of fine-grained spatial cues, limiting performance in complex 3D scenes.
SSM-Based 3D Object Detection. SSM-based detectors represent a new and emerging line of work. LION (Liu et al. 2024) applies a large-group strategy during voxel serialization for more effective spatial context modeling. Voxel Mamba (Zhang et al. 2024b) eliminates the grouping step and serializes the entire voxel space as a single sequence, introducing a Dual-Scale SSM Block to capture hierarchical spatial dependencies. UniMamba combines 3D convolution with SSMs in a compact multi-head architecture that efficiently captures both local and global spatial context.
Although still in their early stages, SSM-based models already outperform traditional CNN-based frameworks on several benchmarks due to their strong modeling capabilities. Our proposed Voxel Diffusion Module (VDM) further enhances these serialized architectures by introducing voxel diffusion and fine-grained spatial feature aggregation prior to serialization. This design supplies Transformer- and SSM-based models with richer input representations, leading to significant performance gains in 3D object detection.
Figure 2: Overall architecture of our proposed method. (a) The Voxel Diffusion Module (VDM) is placed before the sequential detection model, enabling diffusion of raw voxelized point clouds and aggregation of spatial voxel features. Internally, it consists of Submanifold 3D Convolution (SubM3D), Sparse Residual Block (SRB-3D), and Sparse 3D Convolution (SPConv-3D), where SPConv-3D plays a key role in voxel diffusion. (b) Mamba Block. (c) Sparse Residual Block (SRB-3D). The red connection line ”Only-Diffusion” in the figure indicates the model architecture when only the voxel diffusion strategy is applied in the VDM, aiming to highlight the critical impact of voxel diffusion on detection performance.
Method
@@
In this section, we introduce the proposed Voxel Diffusion Module (VDM), designed to enhance the representation of voxelized point cloud features. VDM leverages convolutional operations to perform early-stage voxel diffusion, increasing the density of meaningful voxel information. This provides richer input for downstream serialized models such as Transformers or Mamba. Furthermore, VDM aggregates local spatial features within each voxel’s 3D neighborhood, enabling the sequence models to more effectively utilize spatial positional cues during processing.
Preliminaries
@@
Here, $Q$ (Query), $K$ (Key), and $V$ (Value) represent the input projection vectors. $Q$ interacts with $K^{\top}$ to compute similarity scores, which are then used to weight the values in $V$. The term $d_{k}$ denotes the dimensionality of the key vectors, and the scaling factor $\sqrt{d_{k}}$ is introduced to stabilize the gradients during training. While Transformers are powerful in capturing global dependencies, their quadratic time complexity poses a challenge when handling long sequences, such as in voxelized point clouds.
@@
where $A^{\prime}$ and $B^{\prime}$ are obtained from continuous $A$ and $B$ via a discretization rule: $A^{\prime}=f_{A}(\Delta,A)$ and $B^{\prime}=f_{B}(\Delta,A,B)$. Mamba-based SSM models (Gu and Dao 2023) have demonstrated significant advantages over Transformers in terms of modeling long-range dependencies and computational efficiency. In the context of 3D point cloud detection, such models (Jin et al. 2025; Liu et al. 2024) have shown promising results and even outperformed many Transformer-based architectures.
@@
Sparse Residual Block (SRB). In CNN-based 3D detectors such as SAFDNet (Zhang et al. 2024a), the Sparse Residual Block (SRB) is commonly used for efficient feature extraction. An SRB consists of two 3D submanifold sparse convolution layers and a residual skip connection linking the input to the output. In our design, VDM is composed of stacked SRB modules and additional SubM3D layers to enable both voxel diffusion and spatial feature aggregation. This structure facilitates the enrichment of voxel-level spatial representations before they are passed into the serialized backbone (Transformer or SSM), thereby improving detection accuracy.
@@
Figure 3: Visualization of the Voxel Diffusion Module (VDM) applied to the original point clouds, demonstrating its effect in voxel-level diffusion and feature aggregation.
Overall Architecture
As illustrated in Figure 2, the Voxel Diffusion Module (VDM) is designed to perform voxel-level diffusion of point clouds and aggregate fine-grained voxel features. This facilitates the generation of rich foreground voxel representations and fine-grained spatial features, serving as a beneficial foundation for subsequent transformer-based and state-space model (SSM)-based detectors. Convolutional neural network (CNN)-based models (Zhang et al. 2024a) have been extensively studied in 3D object detection. These models typically begin by applying feature extraction networks to voxelized point clouds, followed by a backbone and a detection head for final predictions. Inspired by this paradigm, the proposed VDM is composed of alternating Submanifold 3D Convolution (SubM3D) layers and Sparse Residual Blocks (SRB). Additionally, two Sparse 3D Convolution (SPConv3D) layers are inserted to downsample and upsample the voxel feature maps, enabling hierarchical feature learning and effective diffusion.
@@
Voxel Diffusion in VDM. Voxel diffusion plays a significant role in enhancing model detection performance, as demonstrated in previous studies (Liu et al. 2024; Zhang et al. 2024a). Inspired by the property of convolution operations to increase the number of sparse features, we introduce a sparse 3D convolution with a kernel size of 3 within our proposed Voxel Diffusion Module (VDM). This module enables diffusion over the input voxel grid, thus expanding the spatial distribution of informative voxels. Figure 3 illustrates how VDM diffuses the original point cloud into a denser voxel representation, providing a richer feature basis for subsequent serial detection networks. To quantitatively evaluate this effect, we perform a statistical analysis on the voxel count post-VDM processing, as shown in Table 1. As observed, the number of total and foreground voxels significantly increases after applying VDM. This expansion ensures that subsequent modules—particularly those based on transformers or State Space Models (SSMs)—can access more comprehensive foreground point features, thereby improving object detection accuracy. It is worth noting that voxel diffusion serves as a crucial component of the VDM. Our later experiments will further highlight its contribution to performance gains, underlining its importance in boosting the effectiveness of serialization-based architectures such as transformers and Mamba variants.
@@
VDM(OD) 216601 10705 74.8
Table 1: Effect of VDM on voxel counts in input point clouds on the Waymo validation dataset. The overall and foreground voxel counts are computed as the total number of corresponding voxels across all frames divided by the total number of frames.
@@
SECOND
CNN
72.3/71.7 63.9/63.3 68.7/58.2 60.7/51.3 60.6/59.3 58.3/57.0 61.0/57.2
PointPillars 72.1/71.5 63.6/63.1 70.6/56.7 62.8/50.3 64.4/62.3 61.9/59.9 62.8/57.8
CenterPoint 74.2/73.6 66.2/65.7 76.6/70.5 68.8/63.2 72.3/71.1 69.7/68.5 68.2/65.8
PV-RCNN 78.0/77.5 69.4/69.0 79.2/73.0 70.4/64.7 71.5/70.3 69.0/67.8 69.6/67.2
PillarNet-18 78.2/77.7 70.4/69.9 79.8/72.6 71.6/64.9 70.4/69.3 67.8/66.7 69.9/67.2
FSD 79.2/78.8 70.5/70.1 82.6/77.3 73.9/69.1 77.1/76.0 74.4/73.3 72.9/70.8
AFDetV2 77.6/77.1 69.7/69.2 80.2/74.6 72.2/67.0 73.7/72.7 71.0/70.1 71.0/68.8
PillarNeXt 78.4/77.9 70.3/69.8 82.5/77.1 74.9/69.8 73.2/72.2 70.6/69.6 71.9/69.7
VoxelNext 78.2/77.7 69.9/69.4 81.5/76.3 73.5/68.6 76.1/74.9 73.3/72.2 72.2/70.1
CenterFormer 75.0/74.4 69.9/69.4 78.6/73.0 73.6/68.3 72.3/71.3 69.8/68.8 71.1/68.9
PV-RCNN++ 79.3/78.8 70.6/70.2 81.3/76.3 73.2/68.0 73.7/72.7 71.2/70.2 71.7/69.5
ConQueR 76.1/75.6 68.7/68.2 79.0/72.3 70.9/64.7 73.9/72.5 71.4/70.1 70.3/67.7
FocalFormer3D –/– 68.1/67.6 –/– 72.7/66.8 –/– 73.7/72.6 71.5/69.0
HEDNet 81.1/80.6 73.2/72.7 84.4/80.0 76.8/72.6 78.7/77.7 75.8/74.9 75.3/73.4
VPF 80.2/79.7 71.9/71.5 82.5/76.9 74.8/69.4 77.1/76.0 74.2/73.2 73.6/71.4
TransFusion
Transformer
–/– –/65.1 –/– –/63.7 –/– –/65.9 –/64.9
SST_TS 76.2/75.8 68.0/67.6 81.4/74.0 72.8/65.9 –/– –/– –/–
SWFormer 77.8/77.3 69.2/68.8 80.9/72.7 72.5/64.9 –/– –/– –/–
FlatFormer - 69.0/68.6 - 71.5/65.3 –/– 68.6/67.5 69.7/67.2
OcTr 78.1/77.6 69.8/69.3 80.8/74.4 72.5/66.5 72.6/71.5 69.9/68.9 70.7/68.2
DSVT-Pillar 79.3/78.8 70.9/70.5 82.8/77.0 75.2/69.8 76.4/75.4 73.6/72.7 73.2/71.0
DSVT-Voxel 79.2/78.8 70.9/70.5 82.9/77.7 75.2/70.3 77.2/76.2 74.3/73.4 73.5/71.4
@@
Voxel Mamba
SSMs
80.8/80.3 72.6/72.2 85.0/80.8 77.7/73.6 78.6/77.6 75.7/74.8 75.3/73.6
LION 80.3/79.9 72.0/71.6 85.8/81.4 78.5/74.3 80.1/79.0 77.2/76.2 75.9/74.0
UniMamba 80.6/80.1 72.3/71.8 86.0/81.3 78.7/74.1 80.3/79.3 77.5/76.5 76.1/74.1

### v3

# Voxel Densification for Serialized 3D Object Detection: Mitigating Sparsity via Pre-serialization Expansion

@@ Abstract
Recent advances in point cloud object detection have increasingly adopted Transformer-based and State Space Models (SSMs) to capture long-range dependencies. However, these serialized frameworks strictly maintain the consistency of input and output voxel dimensions, inherently lacking the capability for voxel expansion. This limitation hinders performance, as expanding the voxel set is known to significantly enhance detection accuracy, particularly for sparse foreground objects. To bridge this gap, we propose a novel Voxel Densification Module (VDM). Unlike standard convolutional stems, VDM is explicitly designed to promote pre-serialization spatial expansion. It leverages sparse 3D convolutions to propagate foreground semantics to neighboring empty voxels, effectively densifying the feature representation before it is flattened into a sequence. Simultaneously, VDM incorporates residual sparse blocks to aggregate fine-grained local context, ensuring rich geometric feature extraction. To balance the computational overhead of increased voxel density, we introduce a strategic cascaded downsampling mechanism. We integrate VDM into both Transformer-based (DSVT) and SSM-based (LION) detectors. Extensive experiments demonstrate that VDM consistently improves detection accuracy across multiple benchmarks. Specifically, our method achieves 74.8 mAPH (L2) on the Waymo validation set and 70.5 mAP on the nuScenes test set. Furthermore, it attains 42.6 mAP on the Argoverse 2 validation set and 67.6 mAP on the ONCE validation set, consistently outperforming the baseline models. The source code will be made publicly available at https://github.com/qifeng22/VDM.
I Introduction
@@
3D object detection based on point clouds plays a crucial role in applications such as autonomous driving [2] and robotic navigation [34]. Inspired by the remarkable success of Transformers [22] and State Space Models (SSMs) [8] in natural language processing, researchers have recently explored adapting these architectures to computer vision tasks, achieving strong performance. For example, in the domain of point cloud 3D object detection, DSVT [24] introduces a dynamic sparse window attention mechanism that adaptively partitions each window into local sub-regions based on voxel sparsity, enabling fully parallel attention computations. LION [14] further extends this direction by employing large grouping windows and leveraging SSMs’ ability to capture long-range voxel dependencies, achieving impressive performance on several public datasets.
Despite these advancements, a fundamental limitation persists in current serialized frameworks: the spatial-serial gap. Both Transformer-based and SSM-based models typically flatten sparse voxels into 1D sequences to model long-range dependencies. However, this serialization process strictly maintains the input-output token consistency, meaning the set of active voxels remains static throughout the backbone. Consequently, these models lack the inherent capability to perform spatial expansion that is trivial in convolutional networks. This limitation is critical because raw point clouds are often sparse and incomplete; without expanding the voxel set to include neighboring empty regions, the model struggles to capture sufficient local context for accurate localization, particularly for distant or occluded objects. While prior works like LION [14] attempt to mitigate this via intermediate feature diffusion, they operate on already-serialized features, where spatial topology is disrupted. In contrast, we argue that the optimal strategy is to explicitly densify the voxel representation before serialization, thereby bridging the gap between spatial completeness and sequential modeling efficiency.
Figure 1: Comparison of serialized detection paradigms. (a) Existing methods such as LION [14] and DSVT [24] directly process sparse voxel sets. (b) Our proposed framework incorporates the Voxel Densification Module (VDM) to explicitly densify and aggregate voxel features before sequence modeling.
Motivated by the spatial flexibility of CNN-based architectures, we propose a novel Voxel Densification Module (VDM), as illustrated in Figure 1, specifically tailored to serve as a pre-serialization stem for Mamba-based and Transformer-based 3D detectors. Positioned prior to the flattening of voxel sequences, VDM leverages sparse 3D convolutional operations to explicitly expand the foreground voxel set, effectively densifying the feature representation. This process generates new informative voxels in the 3D neighborhood of the original sparse points, ensuring that the subsequent serialized model receives a spatially enriched input. Structurally, VDM integrates stacked sparse 3D convolutions with residual blocks to achieve two goals simultaneously: (1) Voxel Densification, which propagates semantics to empty spatial neighbors, and (2) Fine-grained Aggregation, which captures local geometric details. Crucially, to balance the computational cost of processing increased voxel counts, we introduce a strategic downsampling mechanism (reducing resolution to 1/4), optimizing the trade-off between spatial density and sequence length.
@@
VDM is designed to be modular and general, allowing seamless integration with both Transformer-based and SSM-based detectors. To evaluate its effectiveness, we perform extensive experiments on four large-scale datasets. Our VDM-enhanced SSM-based frameworks achieve 74.8 mAPH (L2) on the Waymo validation set, 70.5 mAP on the nuScenes test set, 42.6 mAP on the Argoverse 2 validation set, and 67.6 mAP on the ONCE validation set, consistently outperforming baseline models across different benchmarks. Additionally, the Transformer-based variant (VDM-DSVT) improves performance by +1.2 mAPH (L2) over the DSVT baseline on the Waymo dataset, confirming the general effectiveness of our approach.
@@
  • We identify the limitation of static voxel sets in serialized 3D detectors and propose VDM (Voxel Densification Module), a plug-and-play operator that explicitly densifies sparse voxel inputs via pre-serialization convolution.
  • We design a dual-path architecture within VDM that simultaneously performs spatial expansion (to increase foreground recall) and fine-grained feature aggregation (to enhance local context).
  • VDM demonstrates universal compatibility, seamlessly integrating with both Transformer-based and SSM-based backbones to significantly boost performance.
  • We validate the effectiveness of VDM through extensive experiments, achieving state-of-the-art results across Waymo, nuScenes, Argoverse 2, and ONCE datasets.
II Related Work
3D object detection based on point clouds has seen significant progress. Existing approaches can be broadly categorized into CNN-based, Transformer-based, and State Space Model (SSM)-based methods. While CNN-based models have established a strong baseline, recent serialization-based architectures (Transformers and SSMs) have shown superior potential in capturing long-range dependencies, though they face distinct challenges regarding spatial continuity.
CNN-Based 3D Object Detection. CNN-based detectors are the pioneers in this field, utilizing either voxel-based or point-based representations. VoxelNet [32] introduces Voxel Feature Encoding (VFE) to generate unified feature representations. To improve efficiency, PointPillars [12] collapses point clouds into vertical columns, enabling fast 2D convolution processing. CenterPoint [27] advances this by adopting a center-based anchor-free detection head. More recently, methods like HEDNet [29] utilize encoder-decoder designs for long-range perception. Notably, SAFDNet [28] identifies the issue of feature sparsity and proposes an adaptive feature diffusion strategy. However, SAFDNet primarily focuses on diffusing features within the backbone for CNN-based detectors. In contrast, our work targets the specific ”spatial-serial gap” in serialization-based models, aiming to physically densify the voxel set prior to sequence flattening.
Transformer-Based 3D Object Detection. Transformers have demonstrated state-of-the-art performance by treating voxels as sequences. PVT-SSD [26] employs a Point-Voxel Transformer to capture global dependencies. FlatFormer [15] improves efficiency by flattening point clouds via window-based sorting. DSVT [24] introduces Dynamic Sparse Window Attention to handle sparsity efficiently. Despite their success, these methods typically operate on a static set of non-empty voxels. Due to the strict input-output consistency of attention mechanisms, they lack the inherent ability to expand the voxel set to neighboring empty regions, often leading to insufficient context for sparse objects.
SSM-Based 3D Object Detection. Mamba-based SSMs represent a burgeoning direction. LION [14] leverages SSMs with large window grouping to model wider spatial contexts. Voxel Mamba [30] serializes the entire voxel space to capture global dependencies via Dual-Scale SSM Blocks. UniMamba [11] integrates 3D convolutions with SSMs in a multi-head architecture. While UniMamba combines CNNs and SSMs, it treats them as parallel or interleaved feature extractors.
Different from these approaches, our proposed Voxel Densification Module (VDM) is explicitly designed as a pre-serialization stem. Instead of merely extracting features, VDM fundamentally alters the spatial distribution of the input by densifying foreground voxels. This ensures that the downstream serialized models (whether Transformer or SSM) receive a spatially enriched token sequence, directly addressing the sparsity limitation inherent in serialization.
Figure 2: Overall architecture of our proposed method. (a) The Voxel Densification Module (VDM) is positioned prior to serialization, specifically designed to densify sparse voxel inputs and aggregate local context. The core VDM path (gray box) utilizes stacked Submanifold 3D Convolutions (SubM-3D) and Sparse Residual Blocks (SRB-3D) with downsampling layers (Spconv-3D) to reduce resolution to 1/4. (b) The Mamba Block used in SSM-based backbones. (c) The detailed structure of the Sparse Residual Block (SRB-3D). Note: The red path labeled “Only-Densification” represents the ablation variant where only a single expansion layer (Spconv-3D with $K=3,S=1$) is applied, isolating the efficacy of the voxel expansion mechanism.
III Method
@@
In this section, we introduce the proposed Voxel Densification Module (VDM), designed to enhance the representation of voxelized point cloud features. VDM leverages convolutional operations to perform early-stage voxel expansion, increasing the density of meaningful voxel information. This provides richer input for downstream serialized models such as Transformers or Mamba. Furthermore, VDM aggregates local spatial features within each voxel’s 3D neighborhood, enabling the sequence models to more effectively utilize spatial positional cues during processing.
III-A Preliminaries
@@
Here, $Q$ (Query), $K$ (Key), and $V$ (Value) represent the input projection vectors. While Transformers are powerful in capturing global dependencies, their quadratic time complexity poses a challenge when handling long sequences, such as in voxelized point clouds.
@@
where $A^{\prime}$ and $B^{\prime}$ are obtained from continuous $A$ and $B$ via a discretization rule: $A^{\prime}=f_{A}(\Delta,A)$ and $B^{\prime}=f_{B}(\Delta,A,B)$. Mamba-based SSM models [8] have demonstrated significant advantages over Transformers in terms of modeling long-range dependencies and computational efficiency. In the context of 3D point cloud detection, such models [11, 14] have shown promising results and even outperformed many Transformer-based architectures.
@@
Sparse Residual Block (SRB). In CNN-based 3D detectors such as SAFDNet [28], the Sparse Residual Block (SRB) is commonly used for efficient feature extraction. In our design, VDM is composed of stacked SRB modules and additional SubM3D layers to enable both voxel expansion and spatial feature aggregation. This structure facilitates the enrichment of voxel-level spatial representations before they are passed into the serialized backbone.
III-B Overall Architecture
As illustrated in Figure 2, the Voxel Densification Module (VDM) is designed to perform voxel-level densification of point clouds and aggregate fine-grained voxel features. This facilitates the generation of rich foreground voxel representations, serving as a beneficial foundation for subsequent Transformer-based and SSM-based detectors.
Inspired by CNN-based paradigms, the proposed VDM is composed of alternating Submanifold 3D Convolution (SubM3D) layers and Sparse Residual Blocks (SRB). Additionally, two Sparse 3D Convolution (SPConv3D) layers are inserted to downsample the voxel feature maps, enabling hierarchical feature learning and effective spatial expansion.
@@ III-B Overall Architecture
Figure 3: Visualization of the Voxel Densification Module (VDM) applied to the original point clouds, demonstrating its effect in voxel-level expansion and feature aggregation.
@@
Voxel Densification in VDM. Voxel densification plays a significant role in enhancing model detection performance. Inspired by the property of convolution operations to increase the number of sparse features (dilation), we introduce a sparse 3D convolution with a kernel size of 3 within our proposed VDM. This module enables expansion over the input voxel grid, thus enlarging the spatial distribution of informative voxels. Figure 3 illustrates how VDM expands the original point cloud into a denser voxel representation, providing a richer feature basis for subsequent serial detection networks. To quantitatively evaluate this effect, we perform a statistical analysis on the voxel count post-VDM processing, as shown in Table I. As observed, the number of total and foreground voxels significantly increases after applying VDM. This expansion ensures that subsequent modules—particularly those based on Transformers or SSMs—can access more comprehensive foreground point features, thereby improving object detection accuracy. It is worth noting that voxel expansion serves as a crucial component of the VDM. Our later experiments will further highlight its contribution to performance gains.
@@
Fine-Grained Feature Aggregation. To capture intricate spatial details, we design VDM as a hierarchical framework. As shown in Figure 2(a), the input sparse voxel grid is first processed by a Submanifold 3D Convolution (SubM-3D) stem to generate the initial feature embedding. Let $\mathbf{V}^{(0)}=(\mathcal{P}^{(0)},\mathcal{F}^{(0)})$ denote this initial state. Here, $\mathcal{P}^{(0)}=\{p_{i}\mid p_{i}\in\mathbb{Z}^{3}\}_{i=1}^{N_{0}}$ represents the set of discrete coordinates for $N_{0}$ active voxels, and $\mathcal{F}^{(0)}\in\mathbb{R}^{N_{0}\times C}$ denotes the corresponding feature vectors.
To achieve effective voxel densification and context aggregation, we subsequently employ a cascaded downsampling strategy consisting of $L=2$ stages. Each stage $l$ comprises a sequence of Sparse Residual Blocks (SRB-3D) for feature refinement, followed by a strided Sparse 3D Convolution (SpConv-3D) for spatial expansion. Formally, we define the stage transformation $\mathbf{V}^{(l)}=\mathcal{H}_{l}(\mathbf{V}^{(l-1)})$ for $l\in\{1,2\}$. The operator $\mathcal{H}_{l}$ is defined as the sequential composition of two mapping functions:
First, the feature refinement step utilizes the SRB-3D module, denoted as $\Psi(\cdot)$, to extract high-level semantics while maintaining the input topology:
$\hat{\mathcal{F}}^{(l-1)}=\Psi(\mathcal{F}^{(l-1)})$ (6)
where $\hat{\mathcal{F}}^{(l-1)}$ denotes the refined features at the current scale.
Second, the geometric expansion and aggregation are performed simultaneously by the SpConv-3D layer with learnable weights $\mathcal{W}^{(l)}$ and a stride of $s=2$. This operation transforms both the geometry and features:
$\mathcal{P}^{(l)}=\left\{\left\lfloor\frac{p+k}{2}\right\rfloor\;\middle|\;p\in\mathcal{P}^{(l-1)},\,k\in\mathcal{K}\right\}$ (7)
$\mathcal{F}^{(l)}=\mathcal{W}^{(l)}\ast\hat{\mathcal{F}}^{(l-1)}$ (8)
Here, $\mathcal{K}$ is the kernel support. Notably, despite the reduction in resolution due to the stride, the cardinality of the active voxel set $|\mathcal{P}^{(l)}|$ remains higher than that of a standard baseline at the equivalent voxel size. This phenomenon, attributed to the spatial diffusion effect of the kernel $\mathcal{K}$ in Equation 7, explicitly bridges the spatial-serial gap. The final output of VDM is the densified sparse tensor at the second stage, i.e., $\mathbf{V}^{\prime}=\mathbf{V}^{(2)}=(\mathcal{P}^{(2)},\mathcal{F}^{(2)})$, which provides the enriched representation (at $1/4$ scale) for serialization.
Integration with SSMs and Transformer Modules. The proposed VDM module can be integrated with both Mamba-based and Transformer-based models. By effectively diffusing sparse voxels and aggregating fine-grained point cloud voxel features, VDM further enhances the detection accuracy of the overall model. For the 3D feature maps processed by VDM, we first serialize them into one-dimensional sequences along the $x$ and $y$ directions. As the sequence length increases, Mamba-based networks are able to adopt larger grouping sizes due to their lower computational complexity compared to Transformer-based counterparts. Given the strong performance of LION [14] and DSVT [24] in the field of 3D object detection, we integrate the VDM module into both networks and conduct experiments to validate its effectiveness.
@@
VDM-OD 216601 10705 74.8
TABLE I: Effect of VDM on voxel counts in input point clouds on the Waymo validation dataset. The overall and foreground voxel counts are computed as the total number of corresponding voxels across all frames divided by the total number of frames.
@@
SECOND [25] CNN 72.3/71.7 63.9/63.3 68.7/58.2 60.7/51.3 60.6/59.3 58.3/57.0 61.0/57.2
PointPillars [12] 72.1/71.5 63.6/63.1 70.6/56.7 62.8/50.3 64.4/62.3 61.9/59.9 62.8/57.8
CenterPoint [27] 74.2/73.6 66.2/65.7 76.6/70.5 68.8/63.2 72.3/71.1 69.7/68.5 68.2/65.8
PillarNet-18 [17] 78.2/77.7 70.4/69.9 79.8/72.6 71.6/64.9 70.4/69.3 67.8/66.7 69.9/67.2
FSD [6] 79.2/78.8 70.5/70.1 82.6/77.3 73.9/69.1 77.1/76.0 74.4/73.3 72.9/70.8
AFDetV2 [9] 77.6/77.1 69.7/69.2 80.2/74.6 72.2/67.0 73.7/72.7 71.0/70.1 71.0/68.8
VoxelNext [5] 78.2/77.7 69.9/69.4 81.5/76.3 73.5/68.6 76.1/74.9 73.3/72.2 72.2/70.1
CenterFormer [33] 75.0/74.4 69.9/69.4 78.6/73.0 73.6/68.3 72.3/71.3 69.8/68.8 71.1/68.9
PV-RCNN++ [19] 79.3/78.8 70.6/70.2 81.3/76.3 73.2/68.0 73.7/72.7 71.2/70.2 71.7/69.5
HEDNet [29] 81.1/80.6 73.2/72.7 84.4/80.0 76.8/72.6 78.7/77.7 75.8/74.9 75.3/73.4
VPF [10] 80.2/79.7 71.9/71.5 82.5/76.9 74.8/69.4 77.1/76.0 74.2/73.2 73.6/71.4
TransFusion [1] Transformer –/– –/65.1 –/– –/63.7 –/– –/65.9 –/64.9
FlatFormer [15] - 69.0/68.6 - 71.5/65.3 –/– 68.6/67.5 69.7/67.2
OcTr [31] 78.1/77.6 69.8/69.3 80.8/74.4 72.5/66.5 72.6/71.5 69.9/68.9 70.7/68.2
DSVT-Pillar [24] 79.3/78.8 70.9/70.5 82.8/77.0 75.2/69.8 76.4/75.4 73.6/72.7 73.2/71.0
DSVT-Voxel [24] 79.2/78.8 70.9/70.5 82.9/77.7 75.2/70.3 77.2/76.2 74.3/73.4 73.5/71.4
@@
Voxel Mamba [30] SSMs 80.8/80.3 72.6/72.2 85.0/80.8 77.7/73.6 78.6/77.6 75.7/74.8 75.3/73.6
LION [14] 80.3/79.9 72.0/71.6 85.8/81.4 78.5/74.3 80.1/79.0 77.2/76.2 75.9/74.0
UniMamba [11] 80.6/80.1 72.3/71.8 86.0/81.3 78.7/74.1 80.3/79.3 77.5/76.5 76.1/74.1